40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5430     Entry time: Fri Sep 16 03:22:11 2011
Author: Anamaria 
Type: Update 
Category: LSC 
Subject: More Refl PDs Work and Attempt at DRMI 

Kiwamu, Keiko, Anamaria

I started today with a different input beam, so I had to realign the REFL path again. Then we measured the RF signal out of the 4 REFL PDs and found them to be too low. We increased the power to around 10mA for each diode, and we can see the right modulation frequency on each diode, though REFL165 is way too weak so we might need an RF amplifier on it. We will measure demod board noise tomorrow.

We had an issue with REFL165 not giving the right DC level, low by a factor of 10, even though it was receiving the same optical power as the others. We fifteen-checked clipping and alignment, then pulled it out and measured it on the test stand - found it to be ok. So I uplugged its power cable at the rack and connected it to the AS165 slot. Problem sloved. Not sure what was wrong with the other power slot.

Then we found REFL55 to be clipping on its black glass, we fixed that. But the REFL55 DC power still changes a lot with seemingly not huge motions of the PRM. We'll investigate more tomorrow.

We added a lens in the path to REFL165 because unlike the others it is a 1mm diode. All diodes have about half a turn to a full turn flatness of maximum (on tiny steering mirror).

We set the whitening gain on all four diodes to 21 db.

Not sure if we should set the power to be different on these diodes since their sensitivity is different to RF, and now REFL11 sees huge signal.

We continued the DRMI locking attempt and brought in the SRC, using AS55I to control it. It kind of works/stays locked. We did manage to get MICH and PRC better controlled than last night, but with SRC in the mix, something is wrong. We have to redo f2a filters on SRM and hopefully things will be better after Jenne's suspension work tomorrow. Oplevs not optimized yet either.

We intend to realign POY beam path so we can monitor power in cavities.

ELOG V3.1.3-